# Almost Surely Safe Alignment of Large Language Models at Inference-Time

### v2

@@ Abstract
Even highly capable large language models (LLMs) can produce biased or unsafe responses, and alignment techniques, such as RLHF, aimed at mitigating this issue, are expensive and prone to overfitting as they retrain the LLM. This paper introduces a novel inference-time alignment approach that ensures LLMs generate safe responses almost surely, i.e., with a probability approaching one. We achieve this by framing the safe generation of inference-time responses as a constrained Markov decision process within the LLM’s latent space. Crucially, we augment a safety state that tracks the evolution of safety constraints and enables us to demonstrate formal safety guarantees upon solving the MDP in the latent space. Building on this foundation, we propose InferenceGuard, a practical implementation that safely aligns LLMs without modifying the model weights. Empirically, we demonstrate InferenceGuard effectively balances safety and task performance, outperforming existing inference-time alignment methods in generating safe and aligned responses.
@@ 1 Introduction
LLMs have demonstrated impressive capabilities across a diverse set of tasks, such as summarization (Koh et al., 2022; Stiennon et al., 2020), code generation (Gao et al., 2023; Chen et al., 2021), and embodied robotics (Mower et al., 2024; Kim et al., 2024). However, since those models are primarily trained on vast, unsupervised datasets, their generated responses can often be biased, inaccurate, or harmful (Deshpande et al., 2023; Ganguli et al., 2022; Weidinger et al., 2021; Gehman et al., 2020). As a result, LLMs require alignment with human values and intentions to ensure their outputs are free from controversial content.
The traditional LLM alignment approach involves fine-tuning the model on human-labeled preference data. For instance, works such as (Ouyang et al., 2022; Christiano et al., 2017) use the reinforcement learning from human feedback (RLHF) framework to fine-tune LLMs, constructing a reward model from human feedback and optimizing the model using standard reinforcement learning algorithms like PPO (Schulman et al., 2017). More recent approaches, such as (Tutnov et al., 2025; Yin et al., 2024; Rafailov et al., 2023), bypass reward learning and instead align pre-trained models directly with human preferences.
RLHF alignment can be costly and risks overfitting partly since they modify the LLM’s model weights. To tackle these challenges, alternative approaches adjust the model’s responses directly at inference time while leaving the pre-trained model weights fixed. Several techniques have been proposed for this purpose, such as Best-of-N (Nakano et al., 2021; Stiennon et al., 2020; Touvron et al., 2023), FUDGE (Yang & Klein, 2021), COLD (Qin et al., 2022), CD-Q Mudgal et al. (2023), RE-control Kong et al. (2024), among others. Importantly, those techniques are designed modularly, allowing the alignment module to integrate seamlessly with the pre-trained model. This modularity enables flexible inference-time reconfigurability and quick adaptation to new reward models and datasets. Moreover, it reduces reliance on resource-intensive and often hard-to-stabilize RL processes inherent in the RLHF paradigm.
Despite the successes of inference-time alignment, these methods’ safety aspects have received limited attention so far. While some works have attempted to tackle those issues in inference-time-generated responses, they mainly focus on prompt-based alignment  (Hua et al., 2024; Zhang et al., 2024b), trainable safety classifiers  Niu et al. (2024); Zeng et al. (2024) or protections against adversarial attacks and jailbreaks  (Dong et al., 2024; Guo et al., 2024; Inan et al., 2023). That said, prompt-based methods cannot be guaranteed to consistently produce safe responses, as ensuring safety is heavily reliant on user intervention, requiring extensive engineering and expertise to manage the model’s output effectively. Trainable classifiers focus only on the safety of decoded responses using hidden states or virtual tokens, ignoring task alignment and lacking theoretical guarantees. Moreover, while adversarial robustness is crucial, our work focuses on the key challenge of generating inherently safe responses from the LLM.
@@
In this work, we aim to develop LLMs that generate safe responses at inference time, following a principled approach that guarantees safety almost surely, i.e., with a probability approaching one. To do so, we reformulate the safe generation of inference-time responses as an instance of constrained Markov decision processes (cMDP). We map the cMDP to an unconstrained one through safety state augmentation, bypassing Lagrangian approaches’ limitations that struggle to balance reward maximization and safety feasibility. Focusing on a practical test-time inference algorithm, we adopt a critic-based approach to solve the augmented MDP, eliminating the need for gradients in the LLM. To ensure efficiency, we train our critic in the latent space of the LLM, keeping it small in size and fast during inference. This shift to the latent space complicates the theoretical framework, requiring extensions from (Hernández-Lerma & Muñoz de Ozak, 1992; Sootla et al., 2022). By doing so, we establish, for the first time, that one can guarantee almost sure safety in the original token space.
@@
To leverage this theoretical guarantee in practice, we build upon the augmented MDP framework and introduce two novel implementations for safe inference-time alignment: one that learns a compact critic in the latent space for cases where safety costs can only be queried after the generation of complete responses and another that leverages direct cost queries for efficient inference-time optimization. Finally, we integrate these components into a lookahead algorithm (e.g., Beam Search or Blockwise Decoding (Mudgal et al., 2023)) proposing InferenceGuard. Our experiments tested InferenceGuard starting from safety-aligned and unaligned models. Our results achieved high safety rates—91.04% on Alpaca-7B and 100% on Beaver-7B-v3. Notably, this was accomplished while maintaining a strong balance with rewards, setting new state-of-the-art.
@@ 2 Background
2.1 LLMs as Stochastic Dynamical Systems
LLMs can be viewed as stochastic dynamical systems, where the model’s behavior evolves probabilistically over time, governed by its internal parameters and the inputs it receives. In this perspective, each new token is generated based on the model’s evolving hidden state (Kong et al., 2024; Zimmer et al., 2024). Formally, an LLM transitions as follows: $\left[\mathbf{h}_{t+1},\mathbf{o}_{t+1}\right]^{\mathsf{T}}=f_{\text{LLM}}(\mathbf{h}_{t},\text{y}_{t})$, with $\text{y}_{t}\sim\text{SoftMax}(\mathbf{W}\mathbf{o}_{t})$. Here, $f_{\text{LLM}}(\cdot)$, denotes the aggregation of all decoding layers, $\text{y}_{t}$ a generated token at each time step $t$, and $\mathbf{o}_{t}$ the logits which are linearly mapped by $\mathbf{W}$ to produce a probability distribution over the vocabulary space. Moreover, $\mathbf{h}_{t}$ comprises all key-value pairs accumulated from previous time steps  [^1] Note, $\mathbf{h}_{t}=\left\{\mathbf{K}^{(l)}_{j},\mathbf{V}_{j}^{(l)}\right\}_{l=1}^{L}$ for $j:[1:t]$, i.e., keys and values from all layers accumulated up to the previous time step.. The system evolves until the end-of-sequence (EOS) token is reached.
2.2 Test-Time Alignment of LLMs
Test-time alignment ensures that a pre-trained LLM generates outputs consistent with desired behaviors by solving a MDP, whose initial state is determined by the test prompt. Rewards/costs for alignment can come from various sources, including but not limited to human feedback (Tutnov et al., 2025; Zhong et al., 2024a), environmental feedback from the task or verifiers (Zeng et al., 2023; Yang et al., 2024; Trinh et al., 2024; An et al., 2024; Liang et al., 2024; Mower et al., 2024), or pre-trained reward models (Wang et al., 2024b; Zhang et al., 2024a; Li et al., 2025b).
Several approaches have been developed to address the challenge of test-time alignment. For instance, beam search with rewards (Choo et al., 2022) extends traditional beam search technique by integrating a reward signal to guide LLM decoding at test time. Monte Carlo tree search, on the other hand, takes a more exploratory approach by simulating potential future token sequences to find the path that maximizes the reward function (Zhang et al., 2024a). Best-of-N (BoN) generates multiple candidate sequences and selects the one with the highest reward  (Stiennon et al., 2020; Nakano et al., 2021; Touvron et al., 2023).
We focus on beam search and look-ahead methods for more scalability when developing InferenceGuard.
Figure 1: To aid readability, this figure summarizes the key transitions and notations used throughout the paper.
@@ 3 Safe Test-Time Alignment of LLMs
We frame the problem of safely aligning LLMs at test time as a constrained Markov decision process (cMDP). As noted in Achiam et al. (2017); Sootla et al. (2022) a cMDP is defined as the following tuple: $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{C}_{\text{task}},\mathcal{C}_{\text{safety}},\mathcal{P},\gamma\right\rangle$, with $\mathcal{S}$ and $\mathcal{A}$ denoting the state and action spaces, respectively. The cost function $\mathcal{C}_{\text{task}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ dictates the task’s cost [^2] We define costs as negative rewards, which transforms the problem into an equivalent cost-based MDP., while $\mathcal{C}_{\text{safety}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents a safety cost, which encodes the constraints that the actor must satisfy during inference. The transition model $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ captures the probability of transitioning to a new state given the current state and action. Meanwhile, the discount factor $\gamma\in[0,1)$ trades off immediate versus long-term rewards. The goal of constrained MDPs is to find a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that minimizes the task’s cost while simultaneously satisfying the safety constraints. Given a safety budget d, we write:
$\min_{\pi}$ $\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}$ (1)
s.t. $\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safety}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}\leq d,$
@@ 3.1 Safe Test-Time Alignment as cMDPs
We treat the generation process of safe test-time alignment of LLMs as the solution to a specific cMDP. We introduce our state variable $\mathbf{s}_{t}=\{\mathbf{x},\mathbf{y}_{<t}\}$, which combines the input prompt $\mathbf{x}$ with the tokens (or partial responses) decoded until step $t$. Our policy generates a new token $\text{y}_{t}$ that we treat as an action in the model’s decision-making process. The transition function $\mathcal{P}$ of our MDP is deterministic, where the state $\mathbf{s}_{t}$ is updated by incorporating the generated action $\text{y}_{t}$, i.e., $\mathbf{s}_{t+1}=\mathbf{s}_{t}\oplus\text{y}_{t}=\{\mathbf{x},\mathbf{y}_{\leq t}\}$. We also assume the existence of two cost functions $\mathcal{C}_{\text{task}}$ and $\mathcal{C}_{\text{safety}}$ to assess the correctness and safety of the LLM’s responses. As described in Section 2.2, those functions can originate from various sources, such as human feedback, environmental verifiers, or pre-trained models. While we conduct experiments with these functions being LLMs (see Section 5), our method can be equally applied across various types of task and safety signals.
Regarding the task’s cost, we assume the availability of a function $\mathcal{C}_{\text{task}}$ that evaluates the alignment of the LLM with the given task. This function assigns costs to the partial response based on the input prompt $\mathbf{x}$ such that:
@@
$\min_{\pi}$ $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\overbrace{\{\mathbf{x},\mathbf{y}_{<t}\}}^{\mathbf{s}_{t}},\text{y}_{t})\Big{]}$ (3)
s.t. $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safe}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t})\Big{]}\leq d.$
The above objective aims to minimize the task cost $\mathcal{C}_{\text{task}}$ while ensuring that the safety cost does not exceed a predefined budget $d$. The expectation is taken over the actions (tokens) generated at each timestep.
@@ 3.2 State Augmented Safe Inference-Time Alignment
We could technically use off-the-shelf algorithms to solve Equation 3, such as applying a Lagrangian approach as proposed in Dai et al. (2023). However, there are two main issues with using these standard algorithms. First, they generally require gradients in the model itself—specifically, the LLM—which we want to avoid since our goal is to perform inference-time alignment without retraining the model. Second, these methods rely on a tunable Lagrangian multiplier, making it challenging to maximize rewards while satisfying almost sure constraints optimally.
Instead of using a Lagrangian approach, we take a different direction by augmenting the state space and extending the method proposed by (Sootla et al., 2022) to large language models. In our approach, we augment the state space of the constrained MDP with a “constraint tracker”, effectively transforming the problem into an unconstrained one. This allows us to apply Bellman equations and conduct rigorous proofs with almost sure constraint satisfaction results. However, applying the techniques and proofs from (Sootla et al., 2022) to our test-time setting is not entirely straightforward due to two main challenges: first, the differences in the constrained MDP setting, and second, the process by which we train critics, as we will demonstrate next.
Augmenting the State Space. The following exposition builds on (Sootla et al., 2022), extending their method to address LLM-specific challengers, an area they did not cover. The core idea is to transform the constrained MDP into an unconstrained one by augmenting the state with an additional variable that tracks the remaining budget of the constraint. While doing so, we must ensure that: PI) our augmented state variable tracks the constraints and maintains the Markovian nature of transition dynamics; and PII) our task cost $\mathcal{C}_{\text{task}}$ accounts for this new state representation and is correctly computed w.r.t. the augmented state.
To solve PI, we investigate the evolution of the constraint in Equation 3 and track a scaled-version of the remaining safety budget $\mathbf{\omega}_{t}=d-\sum_{k=1}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})$ that is defined as $\text{z}_{t}=\mathbf{\omega}_{t-1}/\gamma^{t}$. The update of $\text{z}_{t}$ satisfies:
$\text{z}_{t+1}$ $=(\mathbf{\omega}_{t-1}-\gamma^{t}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma^{t+1}$ (4)
$=(\text{z}_{t}-\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma,\ \ \text{with $\text{z}_{0}=d$.}$
Since the dynamics of $\text{z}_{t}$ are Markovian due to the dependence of $\text{z}_{t+1}$ only on $\text{z}_{t}$, $\text{y}_{t}$ and current the state $\{\mathbf{x},\mathbf{y}_{<t}\}$, we can easily augment our original state space with $\text{z}_{t}$, such that $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$. The original dynamics can also be redefined to accommodate for $\tilde{\mathbf{s}}_{t}$:
@@
As noted in (Sootla et al., 2022), this observation holds when the instantaneous costs are nonnegative, ensuring that the accumulated safety cost cannot decrease. In our case, it is natural to assume that the costs are nonnegative for LLMs, as safety violations or misalignments in the output typically incur a penalty, reflecting the negative impact on the model’s performance or ethical standards.

### v3

# On Almost Surely Safe Alignment of Large Language Models at Inference-Time

@@ Abstract
We introduce a novel inference-time alignment approach for LLMs that aims to generate safe responses almost surely, i.e., with probability approaching one. Our approach models the generation of safe responses as a constrained Markov Decision Process (MDP) within the LLM’s latent space. We augment a safety state that tracks the evolution of safety constraints and dynamically penalize unsafe generations to ensure the generation of safe responses. Consequently, we demonstrate formal safety guarantees w.r.t. the given cost model upon solving the MDP in the latent space with sufficiently large penalties. Building on this foundation, we propose InferenceGuard, a practical implementation that safely aligns LLMs without modifying the model weights. Empirically, we demonstrate that InferenceGuard effectively balances safety and task performance, outperforming existing inference-time alignment methods in generating safe and aligned responses. Our findings contribute to the advancement of safer LLM deployment through alignment at inference-time, thus presenting a promising alternative to resource-intensive, overfitting-prone alignment techniques like RLHF.
@@ 1 Introduction
LLMs have demonstrated impressive capabilities across a diverse set of tasks, such as summarization [1, 2], code generation [3, 4], and embodied robotics [5, 6]. However, since those models are primarily trained on vast, unsupervised datasets, their responses can often be biased, inaccurate, or harmful [7, 8, 9, 10]. To prevent such controversial content, LLMs require alignment with better human values.
The predominant approach for LLM alignment is Reinforcement Learning from Human Feedback (RLHF) [11, 12], which fine-tunes the model using human preference data. A drawback of RLHF, however, is its potential training cost and risk of overfitting, partly because this method modifies the model’s weights. In contrast, inference-time alignment adjusts the model’s outputs directly during inference to align with a reward model, while leaving the model weights fixed ([13, 2, 14]). Despite the successes of inference-time alignment, their safety aspects have received limited attention so far.
@@
In this work, we aim to develop a principled inference-time alignment technique that guarantees the safety of LLM responses almost surely, i.e., with a probability approaching one. To do so, we reformulate the safe generation of inference-time responses as an instance of constrained Markov decision processes (cMDP) with the objective of maximizing task performance while satisfying safety cost constraints. We map the cMDP to an unconstrained one through safety state augmentation, bypassing Lagrangian approaches’ limitations that struggle to balance reward maximization and safety feasibility. Specifically, the tracking of the safety state enables one to dynamically penalize the task reward when constraint violations occur. Focusing on practical efficiency, we adopt a critic-based approach to solve the augmented MDP, eliminating the need for gradients in the LLM. To ensure efficiency, we train our critic in the LLM’s latent space, keeping it small in size and fast during inference. This shift to the latent space complicates the theoretical framework, requiring extensions from previous works [15, 16]. By doing so, we establish, for the first time, that for sufficiently large penalties one can guarantee almost sure safety in the original token space w.r.t. given cost model.
Figure 1: Overview of the InferenceGuard framework. Given a prompt $\mathbf{x}$, InferenceGuard sequentially generates beams of tokens from the base LLM, augments them with the safety state to track the evolution of safety constraints, evaluates each beam using our learned value models for both safety and task alignment, and filters the top $K$ beams (see Section 5.2). If all beams are unsafe, it penalizes the logits of unsafe tokens and resamples. Consequently, InferenceGuard efficiently generates responses with superior safety rates and strong task performance.
@@
To leverage this theoretical guarantee in practice, we build upon the augmented MDP framework and introduce two novel implementations for safe inference-time alignment: one that learns a compact critic in the latent space for cases where safety costs can only be queried after the generation of complete responses and another that leverages direct cost queries for efficient inference-time optimization. Finally, we integrate these components into a lookahead algorithm (e.g., Beam Search or Blockwise Decoding [14]) proposing InferenceGuard. While test-time alignment often introduces additional latency – commonly referred to as the alignment tax – InferenceGuard matches the decoding cost of standard beam search yet achieves markedly high safety rates – 94.46% on Alpaca-7B, 98.45% on Llama3.1-8B-Instruct, 98.97% on Vicuna-7B and 100% on Beaver-7B-v3. Notably, this was accomplished while maintaining a strong balance with rewards, setting new state-of-the-art.
2 Related Work
LLM alignment and safety: Pre-trained LLMs are often aligned to specific tasks using RLHF, where LLMs are fine-tuned with a learned reward model [12, 2, 17, 11] constructed from human feedback using standard reinforcement learning algorithms like PPO [18]. More recent approaches, such as [19, 20, 21, 22, 23, 24, 25, 26], bypass reward learning and instead align pre-trained models directly with human preferences. [27, 8] first applied fine-tuning in the context of safety, and [28] proposed safe fine-tuning via Lagrangian optimization. Other safety-focused methods such as [29, 30, 31, 32, 33, 34, 35, 36, 37] are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety w.r.t. the given cost model during inference.
Inference time alignment: In order to reduce reliance on resource-intensive and often hard-to-stabilize RL processes inherent in the RLHF paradigm, inference-time alignment techniques such as Best-of-N [13, 2, 38, 39], guided decoding (steering token generation based on reward or a trained value function) [40, 41, 14, 42, 43, 44, 45, 46], among others, have been proposed. Importantly, these techniques are designed modularly, allowing the alignment module to integrate seamlessly with the pre-trained model by adjusting the model’s responses directly at inference time. This modularity enables flexible inference-time reconfigurability and quick adaptation to new reward models and datasets.
While a few works have attempted to tackle the safety issue in inference-time-aligned responses, they mainly focus on prompt-based alignment [32, 33, 47, 48], trainable safety classifiers [49, 50] or protections against adversarial attacks and jailbreaks [51, 34, 52, 53]. That said, prompt-based methods cannot be guaranteed to consistently produce safe responses, as ensuring safety is heavily reliant on user intervention, requiring extensive engineering and expertise to manage the model’s output effectively. Trainable classifiers focus only on the safety of decoded responses using hidden states or virtual tokens, ignoring task alignment and lacking theoretical guarantees. Moreover, while adversarial robustness is crucial, our work focuses on the key challenge of generating inherently safe responses from the LLM. Compared to those methods, we are the first to theoretically guarantee almost surely safe alignment w.r.t. given cost model with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints without introducing extra parameters, e.g., Lagrangian multipliers.
We detail other related works extensively in Appendix A.
@@ 3 Background
LLMs can be viewed as stochastic dynamical systems, where the model’s behavior evolves probabilistically over time, governed by its internal parameters and the inputs it receives. In this perspective, each new token is generated based on the model’s evolving hidden state [42, 54]. Formally, an LLM transitions as follows: $\left[\mathbf{h}_{t+1},\mathbf{o}_{t+1}\right]^{\mathsf{T}}=f_{\text{LLM}}(\mathbf{h}_{t},\text{y}_{t})$, with $\text{y}_{t}\sim\text{SoftMax}(\mathbf{W}\mathbf{o}_{t})$. Here, $f_{\text{LLM}}(\cdot)$, denotes the aggregation of all decoding layers, $\text{y}_{t}$ a generated token at each time step $t$, and $\mathbf{o}_{t}$ the logits which are linearly mapped by $\mathbf{W}$ to produce a probability distribution over the vocabulary space. Moreover, $\mathbf{h}_{t}$ comprises all key-value pairs accumulated from previous time steps  [^1] Note, $\mathbf{h}_{t}=\left\{\mathbf{K}^{(l)}_{j},\mathbf{V}_{j}^{(l)}\right\}_{l=1}^{L}$ for $j\in[1,t]$, i.e., keys and values from all layers till time step $t$.. The system evolves until the end-of-sequence (EOS) token is reached.
Test-Time Alignment of LLMs. A pre-trained LLM can be ensured to generate outputs consistent with desired behaviors by solving an MDP, whose initial state is determined by the test prompt. Rewards/costs for alignment can come from various sources, such as human feedback [19, 55], environmental feedback from the task or verifiers [56, 57, 58, 59, 60, 5], or pre-trained reward models [61, 62, 63].
Several approaches address the challenge of test-time alignment. For instance, beam search with rewards [64] extends traditional beam search by integrating a reward signal to guide LLM decoding at test time. Monte Carlo tree search, on the other hand, takes a more exploratory approach by simulating potential future token sequences to find the path that maximizes the reward function [62]. Best-of-N (BoN) generates multiple candidate sequences and selects the one with the highest reward  [2, 13, 38]. We focus on beam search methods for more scalability when developing InferenceGuard.
@@ 4 Safe Test-Time Alignment of LLMs
We frame the problem of safely aligning LLMs at test time as a constrained Markov decision process (cMDP). As noted in [65, 16] a cMDP is defined as the following tuple: $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{C}_{\text{task}},\mathcal{C}_{\text{safety}},\mathcal{P},\gamma\right\rangle$, with $\mathcal{S}$ and $\mathcal{A}$ denoting the state and action spaces, respectively. The cost function $\mathcal{C}_{\text{task}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ dictates the task’s cost [^2] We define costs as negative rewards, which transforms the problem into an equivalent cost-based MDP., while $\mathcal{C}_{\text{safety}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents a safety cost, which encodes the constraints that the actor must satisfy during inference. The transition model $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ captures the probability of transitioning to a new state given the current state and action. Meanwhile, the discount factor $\gamma\in[0,1)$ trades off immediate versus long-term rewards. The goal of constrained MDPs is to find a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that minimizes the task’s cost while simultaneously satisfying the safety constraints. Given a safety budget d, we write:
$\min_{\pi}$ $\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}\quad\text{s.t.}\quad\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safety}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}\leq d,$ (1)
@@ 4.1 Safe Test-Time Alignment as cMDPs
We treat the generation process of safe test-time alignment of LLMs as the solution to a specific cMDP. We introduce our state variable $\mathbf{s}_{t}=\{\mathbf{x},\mathbf{y}_{<t}\}$, which combines the input prompt $\mathbf{x}$ with the tokens (or partial responses) decoded until step $t$. Our policy generates a new token $\text{y}_{t}$ that we treat as an action in the model’s decision-making process. The transition function $\mathcal{P}$ of our MDP is deterministic, where the state $\mathbf{s}_{t}$ is updated by incorporating the generated action $\text{y}_{t}$, i.e., $\mathbf{s}_{t+1}=\mathbf{s}_{t}\oplus\text{y}_{t}=\{\mathbf{x},\mathbf{y}_{\leq t}\}$. We also assume the existence of two cost functions $\mathcal{C}_{\text{task}}$ and $\mathcal{C}_{\text{safety}}$ to assess the correctness and safety of the LLM’s responses. As described in Section 3, those functions can originate from various sources, such as human feedback, environmental verifiers, or pre-trained models. While we conduct experiments with these functions being LLMs (see Section 6), our method can be equally applied across various types of task and safety signals.
We assume the availability of a function $\mathcal{C}_{\text{task}}$ that evaluates the alignment of the LLM with the given task. This function assigns costs to the partial response based on the input prompt $\mathbf{x}$ such that:
@@
$\min_{\pi}$ $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\overbrace{\{\mathbf{x},\mathbf{y}_{<t}\}}^{\mathbf{s}_{t}},\text{y}_{t})\Big{]}\quad\text{s.t.}\quad\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safe}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t})\Big{]}\leq d.$ (3)
The above objective aims to minimize the task cost $\mathcal{C}_{\text{task}}$ while ensuring that the safety cost does not exceed a predefined budget $d$. The expectation is taken over the actions (tokens) generated at each step.
@@ 4.2 State Augmented Safe Inference-Time Alignment
We could technically use off-the-shelf algorithms to solve Equation 3, such as applying a Lagrangian approach as proposed in [28]. However, there are two main issues with using these standard algorithms. First, they generally require gradients in the model itself—specifically, the LLM—which we want to avoid since our goal is to perform inference-time alignment without retraining the model. Second, these methods rely on a tunable Lagrangian multiplier, making it challenging to maximize rewards while satisfying almost sure constraints optimally.
Instead of a Lagrangian approach, we take a different direction by augmenting the state space and extending the method proposed by [16] to large language models. In our approach, we augment the state space of the constrained MDP with a “constraint tracker”, effectively transforming the problem into an unconstrained one. This allows us to apply Bellman equations and conduct rigorous proofs with almost sure constraint satisfaction results. However, applying the techniques and proofs from [16] to our test-time setting is not entirely straightforward due to two main challenges: first, the differences in the constrained MDP setting, and second, the process by which we train critics, as we will demonstrate next.
Augmenting the State Space. The following exposition builds on [16], extending their method to address LLM-specific challengers, an area they did not cover. The core idea is to transform the constrained MDP into an unconstrained one by augmenting the state with an additional variable that tracks the remaining budget of the constraint. While doing so, we must ensure that: PI) our augmented state variable tracks the constraints and maintains the Markovian nature of transition dynamics; and PII) our task cost $\mathcal{C}_{\text{task}}$ accounts for this new state representation.
We solve PI by tracking a scaled-version of the remaining safety budget (see Equation 3 and) $\mathbf{\omega}_{t}=d-\sum_{k=1}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})$, defined as $\text{z}_{t}=\mathbf{\omega}_{t-1}/\gamma^{t}$. The update of $\text{z}_{t}$ satisfies:
$\text{z}_{t+1}=(\mathbf{\omega}_{t-1}-\gamma^{t}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma^{t+1}=(\text{z}_{t}-\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma,\ \ \text{with $\text{z}_{0}=d$.}$ (4)
The dynamics of $\text{z}_{t}$ are Markovian and dependent only on $\text{z}_{t-1}$, $\text{y}_{t-1}$ and current state $\{\mathbf{x},\mathbf{y}_{<t-1}\}$. Hence, we can easily augment our original state space with $\text{z}_{t}$, such that $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$. The original dynamics can also be redefined to accommodate for $\tilde{\mathbf{s}}_{t}$:
@@
As noted in [16], this observation holds when the instantaneous costs are nonnegative, ensuring that the accumulated safety cost cannot decrease. In our case, it is natural to assume that the costs are nonnegative for LLMs, as safety violations or misalignments in the output typically incur a penalty, reflecting the negative impact on the model’s performance or ethical standards.